# Correcting the Mythos of KL-Regularization: Direct Alignment without Overoptimization via $\chi^{2}$-Preference Optimization

### v2

@@ 3.1 Framework: $\boldsymbol{\chi^{2}}$-Regularized Reward Optimization
$\chi^{2}$-divergence is a more aggressive form of regularization than KL-divergence; we have $D_{\mathsf{KL}}\left(\mathbb{P}\,\|\,\mathbb{Q}\right)\leq{}2D_{\chi^{2}}\left(\mathbb{P}\;\|\;{}\mathbb{Q}\right)$, but the converse is not true in general. We consider the following $\chi^{2}$-regularized RL objective:333 We write $D_{\chi^{2}}\left(\pi\;\|\;{}\pi_{\mathsf{ref}}\right)=\operatorname{\mathbb{E}}\big{[}D_{\chi^{2}}\left(\pi(\cdot\mid{}x)\;\|\;{}\pi_{\mathsf{ref}}(\cdot\mid{}x)\right)\big{]}=\operatorname{\mathbb{E}}_{\pi}\left[\frac{\pi(a\mid{}x)}{\pi_{\mathsf{ref}}(a\mid{}x)}\right]$.
$J_{\beta}^{\chi}(\pi)\vcentcolon={}\operatorname{\mathbb{E}}_{\pi}\left[r^{\star}(x,a)\right]-\beta\cdot D_{\chi^{2}}\left(\pi\;\|\;{}\pi_{\mathsf{ref}}\right).$ (6)
Moving to a form of regularization that penalizes deviations from $\pi_{\mathsf{ref}}$ more forcefully than KL-regularization is a natural approach to mitigating overoptimization, but an immediate concern is that this may lead to overly conservative algorithms. As we will show, however, $\chi^{2}$-divergence is better suited to the geometry of offline alignment, as it has the unique property (not shared by KL-divergence) that its value quantifies the extent to which the accuracy of a reward model ${\widehat{r}}$ trained under $\pi_{\mathsf{ref}}$ will transfer to a downstream policy $\pi$ of interest (Lemma E.3). This implies that the $\chi^{2}$-regularized RL objective in Eq. 6 meaningfully implements a form of pessimism in the face of uncertainty, and by tuning the regularization parameter $\beta>0$, we can keep the learned policy $\widehat{\pi}$ close to $\pi_{\mathsf{ref}}$ in the “right” (uncertainty-aware) way and fully utilize the offline dataset $\mathcal{D}_{\mathsf{pref}}$. As such, we view optimizing $\chi^{2}$-regularized rewards, i.e., $\mathop{\mathrm{argmax}}_{\pi\in\Pi}J_{\beta}^{\chi}(\pi)$ as a general principle to guide algorithm design for offline alignment (as well as offline RL more broadly), which we expect to find broader use.
@@
While this indeed leads to meaningful statistical guarantees (cf. Appendix C), we adopt a simpler and more direct approach inspired by DPO, which removes the need for a separate reward estimation step.
@@ 3.2 The $\boldsymbol{\chi}$PO Algorithm
In other words, we use both $\chi^{2}$-regularization and KL-regularization; $\chi^{2}$-regularization enforces pessimism, while KL-regularization enforces the barrier property and facilitates reparameterization. Indeed, the link function $\phi$ (Eq. 8) used in $\chi$PO has $\phi(z)\vcentcolon={}f_{\chi_{\textsf{mix}}}^{\prime}(z)=z+\log{}z$, which satisfies $0\notin{}\mathrm{dom}(f_{\chi_{\textsf{mix}}}^{\prime})$, so Eq. 10 yields the reparameterization $r^{\star}(x,a)=\beta\phi\left(\frac{\pi^{\star}_{\beta;f_{\chi_{\textsf{mix}}}}(a\mid{}x)}{\pi_{\mathsf{ref}}(a\mid{}x)}\right)+Z_{\beta,r^{\star};f_{\chi_{\textsf{mix}}}}(x)$. Substituting this identity into the maximum likelihood estimation objective (Eq. 3) yields the $\chi$PO algorithm.
@@ 3.3 Theoretical Guarantees
Our second assumption asserts that the “implicit” reward models induced by the policy class $\Pi$ in $\chi$PO have bounded range.
@@
Finally, to quantify the coverage of the offline preference dataset $\mathcal{D}_{\mathsf{pref}}$ generated by $\pi_{\mathsf{ref}}$, we make use of $L_{1}$-concentrability (Farahmand et al., 2010; Xie and Jiang, 2020; Zanette et al., 2021), a sharper notion of coverage than the $L_{\infty}$-concentrability coefficient discussed in the prequel.
Definition 3.1 ($L_{1}$-Concentrability).
The single-policy $L_{1}$-concentrability coefficient for a policy $\pi$ is given by
$\mathcal{C}^{\pi}\vcentcolon={}\operatorname{\mathbb{E}}_{\pi}\left[\frac{\pi(a\mid x)}{\pi_{\mathsf{ref}}(a\mid x)}\right].$
For a given policy $\pi$, $\mathcal{C}^{\pi}$ measures the extent to which $\pi$ is covered by $\pi_{\mathsf{ref}}$. Note that $L_{1}$-concentrability is equivalent to $\chi^{2}$-divergence up to constant shift ($\mathcal{C}^{\pi}=1+2D_{\chi^{2}}\left(\pi\;\|\;{}\pi_{\mathsf{ref}}\right)$), and that in general, we can have $\mathcal{C}^{\pi}\ll\mathcal{C}_{\infty}^{\pi}$.
The following is our main sample complexity guarantee for $\chi$PO.
@@
Theorem 3.1 shows that $\chi$PO achieves a sample complexity guarantee that scales only with the single-policy concentrability parameter $\mathcal{C}^{\pi^{\star}}$ for the comparator policy $\pi^{\star}$, for all policies $\pi^{\star}$ simultaneously. In particular, roughly $n=O\left(\frac{\mathcal{C}^{\pi^{\star}}\log(\lvert\Pi\rvert/\delta)}{\varepsilon^{2}}\right)$ examples are sufficient to learn a policy that is $\varepsilon$-suboptimal relative to $\pi^{\star}$. The dependence on single-policy concentrability in Theorem 3.1 reflects the algorithm’s robustness to overoptimization. In particular, this means that the algorithm is competitive with any policy $\pi^{\star}$ that is covered by the offline dataset $\mathcal{D}_{\mathsf{pref}}$ (in the sense that $\mathcal{C}^{\pi^{\star}}=O(1)$), which is effectively the best one can hope for in a purely offline setting. In contrast, naive offline alignment methods like DPO have sample complexity that scales with all-policy concentrability (roughly, $\max_{\pi}\mathcal{C}^{\pi}$), even when the comparator policy $\pi^{\star}$ is sufficiently covered (Zhu et al., 2023; Song et al., 2024). To highlight this, in Section 4 we give a concrete example in which $\chi$PO allows the user to tune $\beta$ to achieve tight statistical rates, yet no choice of $\beta$ for DPO leads to comparable performance (effectively, any choice of $\beta$ for DPO is either susceptible to overoptimization, or has unacceptably high bias). All prior works that achieve similar sample complexity guarantees based on single-policy concentrability are either impractical, or require more restrictive statistical assumptions on the policy class (Ye et al., 2024; Liu et al., 2024; Cen et al., 2024; Fisch et al., 2024; Ji et al., 2024). [^4] A notable difference is that some of these works achieve guarantees based on tighter coverage parameters than single-policy $L_{1}$-concentrability, which reflect the structure of the policy or reward class. This appears to be out of reach for our techniques.
Regarding the parameter $V_{\mathsf{max}}$, we observe that since the policy $\pi^{\star}_{{\scriptscriptstyle\beta}}$ satisfies $\left\lvert\beta\phi\left(\frac{\pi^{\star}_{{\scriptscriptstyle\beta}}(a\mid{}x)}{\pi_{\mathsf{ref}}(a\mid{}x)}\right)-\beta\phi\left(\frac{\pi^{\star}_{{\scriptscriptstyle\beta}}(b\mid{}x)}{\pi_{\mathsf{ref}}(b\mid{}x)}\right)\right\rvert\leq 2R_{\mathsf{max}}$, information-theoretically we can always achieve $V_{\mathsf{max}}=2R_{\mathsf{max}}$ by pre-filtering the policy class $\Pi$ to remove all policies for which this inequality does not hold. Since this may be non-trivial in practice, we incorporate clipping in Eq. 9, which, for precisely the above reason, we expect to improve performance empirically. In Section 4, we discuss the role of the $V_{\mathsf{max}}$ parameter and Assumption 3.2 in greater depth. See also guarantees for the $\chi^{2}$-RLHF algorithm in Appendix C, which avoid dependence on this parameter.
@@
Consider the setting in Example 3.1, where the policy class $\Pi_{\mathcal{R},\beta}$ is the set of mixed $\chi^{2}$-regularized policies induced by a reward model class $\mathcal{R}$ with $r^{\star}\in\mathcal{R}$ and $\beta>0$. For any $\delta\in(0,1)$, there exists a choice [^5] It is unclear how to select $\beta^{\star}$ in a data-driven manner as it depends non-trivially on the functionals $\pi\mapsto C^{\pi}$ and $\pi\mapsto J(\pi)$. for $\beta^{\star}>0$ such that with probability at least $1-\delta$, $\chi$PO (Algorithm 1), with class $\Pi_{\mathcal{R},\beta^{\star}}$, produces a policy $\widehat{\pi}$ such that for all policies $\pi^{\star}$ simultaneously, we have
@@
Specializing to the case of multi-armed bandits, we believe that the sample complexity bound in Eq. 13 is optimal in general (Rashidinejad et al., 2021). Note that while we consider finite classes $\Pi$ in Theorem 3.1 for simplicity, extension to infinite classes is trivial via standard uniform convergence arguments. We remark that the exponential dependence on $R_{\mathsf{max}}$ in Theorem 3.1 is an intrinsic feature of the Bradley-Terry model, and can be found in all prior work (Rosset et al., 2024; Xie et al., 2024). Finally, we remark that the weight on the KL term in the mixed $\chi^{2}$-regularized objective is not important for our statistical guarantees. For any $\gamma\in(0,1]$, we can replace the link function $\phi(\cdot)$ in $\chi$PO with
@@
which corresponds to the regularized objective $J_{\beta,\gamma}^{\chi_{\textsf{mix}}}(\pi)=\operatorname{\mathbb{E}}_{\pi}\left[r^{\star}(x,a)\right]-\beta\cdot D_{\chi^{2}}\left(\pi\;\|\;{}\pi_{\mathsf{ref}}\right)-\gamma\beta\cdot D_{\mathsf{KL}}\left(\pi\,\|\,\pi_{\mathsf{ref}}\right)$. This leads to identical guarantees for any $\gamma\in(0,1]$ (Theorem E.1 in Appendix E); essentially, we only require that $\gamma$ is positive to ensure the reparameterization in Eq. 10 is admissible.
@@ 4.2 The Bias-Overoptimization Tradeoff
Bias. Compared to DPO, $\chi$PO has a higher probability of taking the optimal action $a_{\mathsf{good}}$, and thus has lower bias (Center). As a result, it strikes a better better bias-overoptimization tradeoff than DPO, and is competitive with respect to the comparator $a_{0}$ even when DPO fails to converge.
@@ 4.3 An Illustrative Example
Figure 3: The regret $J(a_{0})-J(\widehat{\pi})$ in the construction from Proposition A.1 for different values of $n$. We again condition on the “bad” event $\mathcal{E}$ where ${\widehat{r}}=r_{2}\neq r^{\star}$. For each $n$, the error from overoptimization dominates when $\beta\leq(2\log n)^{-1}$ (as discussed in Section 4.3), and the error from bias dominates when $\beta>(2\log n)^{-1}$. Taking the best choice of $\beta$ for each method, DPO converges at an exponentially slower rate than $\chi$PO (see Proposition A.1 for formal statement).
@@ 4.4 Nontriviality and Role of $V_{\mathsf{max}}$ Parameter
Assumption 3.2 effectively implies that all policies $\pi\in\Pi$ satisfy $\big{\|}\frac{\pi}{\pi_{\mathsf{ref}}}\big{\|}_{\infty}\lesssim\frac{V_{\mathsf{max}}}{\beta}$; in other words, the policy class we use in $\chi$PO satisfies all-policy $L_{\infty}$-concentrability with $\max_{\pi\in\Pi}\mathcal{C}^{\pi}_{\infty}\lesssim\frac{V_{\mathsf{max}}}{\beta}$. This might seem to trivialize the offline alignment problem at first glance, since it would suffice to prove a generalization guarantee based on all-policy concentrability, and then plug this bound in. We will show that this is not the case, and that this is actually an intrinsic feature of $\chi^{2}$-regularization.
@@
In this section, we sketch the proof of the main guarantee for $\chi$PO, Theorem 3.1, with the full proof deferred to Appendix E. A central object in the proof is the implicit reward model induced by the $\chi$PO policy $\widehat{\pi}$, which we define via
@@
As we will show, this reward model is a natural bridge between $\chi$PO and the corresponding mixed $\chi^{2}$-regularized RLHF objective in Section 3.1, and allows us to view $\chi$PO from a reward-based perspective. In particular, note that if we analogously define an induced reward model class $\mathcal{R}_{\Pi}\vcentcolon={}\{r(x,a)=\beta\phi{}\left(\frac{\pi(a\mid{}x)}{\pi_{\mathsf{ref}}(a\mid{}x)}\right):\pi\in\Pi\}$, then 3 of $\chi$PO can be viewed as performing maximum likelihood estimation over this class (in the sense of Eq. 3) under the Bradley-Terry model. Under Assumption 3.1, $\mathcal{R}_{\Pi}$ realizes the true reward function $r$ up to an action-independent shift. As a result, if we define $\Delta^{{r}}(x,a,b)\vcentcolon={}r(x,a)-r(x,b)$, then using a fairly standard generalization bound for maximum likelihood estimation (e.g., Wong and Shen (1995); Zhang (2006); de Geer (2000); see Lemma E.1), we can show that
@@
Lemma 5.1 (Informal version of Lemma E.3).
@@
Lemma 5.2 (Informal version of Lemma E.2).
@@
6 General Preference Models

### v3

@@ 3.1 Framework: $\boldsymbol{\chi^{2}}$-Regularized Reward Optimization
$\chi^{2}$-divergence is a more aggressive form of regularization than KL-divergence; we have $D_{\mathsf{KL}}\left(\mathbb{P}\,\|\,\mathbb{Q}\right)\leq{}2D_{\chi^{2}}\left(\mathbb{P}\;\|\;{}\mathbb{Q}\right)$, but the converse is not true in general. We consider the following $\chi^{2}$-regularized RL objective:333 Note the definition of $D_{\chi^{2}}\left(\pi\;\|\;{}\pi_{\mathsf{ref}}\right)$ differs from $\operatorname{\mathbb{E}}[D_{\chi^{2}}\left(\pi(\cdot\mid{}x)\;\|\;{}\pi_{\mathsf{ref}}(\cdot\mid{}x)\right)]$ only by a constant scaling and shift, both of which are inconsequential when used as regularization in an optimization objective.
$J_{\beta}^{\chi}(\pi)\vcentcolon={}\operatorname{\mathbb{E}}_{\pi}\left[r^{\star}(x,a)\right]-\beta\cdot D_{\chi^{2}}\left(\pi\;\|\;{}\pi_{\mathsf{ref}}\right),\qquad D_{\chi^{2}}\left(\pi\;\|\;{}\pi_{\mathsf{ref}}\right):=\operatorname{\mathbb{E}}_{\pi}\left[\frac{\pi(a\mid{}x)}{\pi_{\mathsf{ref}}(a\mid{}x)}\right].$ (6)
Moving to a form of regularization that penalizes deviations from $\pi_{\mathsf{ref}}$ more forcefully than KL-regularization is a natural approach to mitigating overoptimization, but an immediate concern is that this may lead to overly conservative algorithms. As we will show, however, $\chi^{2}$-divergence is better suited to the geometry of offline alignment, as it has the unique property (not shared by KL-divergence) that its value quantifies the extent to which the accuracy of a reward model ${\widehat{r}}$ trained under $\pi_{\mathsf{ref}}$ will transfer to a downstream policy $\pi$ of interest (Lemma F.3). This implies that the $\chi^{2}$-regularized RL objective in Eq. 6 meaningfully implements a form of pessimism in the face of uncertainty, and by tuning the regularization parameter $\beta>0$, we can keep the learned policy $\widehat{\pi}$ close to $\pi_{\mathsf{ref}}$ in the “right” (uncertainty-aware) way. As such, we view optimizing $\chi^{2}$-regularized rewards, i.e., $\mathop{\mathrm{argmax}}_{\pi\in\Pi}J_{\beta}^{\chi}(\pi)$ as a general principle to guide algorithm design for offline alignment (as well as offline RL more broadly), which we expect to find broader use.
@@
While this indeed leads to strong statistical guarantees (cf. Appendix B), we adopt a simpler and more direct approach inspired by DPO, which removes the need for a separate reward estimation step.
@@ 3.2 The $\boldsymbol{\chi}$PO Algorithm
In other words, we use both $\chi^{2}$-regularization and KL-regularization; $\chi^{2}$-regularization enforces pessimism, while KL-regularization enforces the barrier property and facilitates reparameterization. Indeed, the link function $\phi$ (Eq. 8) used in $\chi$PO has $\phi(z)\vcentcolon={}f_{\chi_{\textsf{mix}}}^{\prime}(z)=z+\log{}z$, which satisfies $0\notin{}\mathrm{dom}(f_{\chi_{\textsf{mix}}}^{\prime})$, so Eq. 10 yields the reparameterization $r^{\star}(x,a)=\beta\phi\Big{(}\frac{\pi^{\star}_{\beta;f_{\chi_{\textsf{mix}}}}(a\mid{}x)}{\pi_{\mathsf{ref}}(a\mid{}x)}\Big{)}+Z_{\beta,r^{\star};f_{\chi_{\textsf{mix}}}}(x)$. Substituting this identity into the maximum likelihood estimation objective (Eq. 3) yields the $\chi$PO algorithm.
@@ 3.3 Theoretical Guarantees
Our second assumption asserts that the implicit reward models induced by the policy class $\Pi$ in $\chi$PO have bounded range.
@@
[…]
@@
Theorem 3.1 shows that $\chi$PO achieves a sample complexity guarantee that scales only with the single-policy concentrability parameter $\mathcal{C}^{\pi^{\star}}$ for the comparator policy $\pi^{\star}$, for all policies $\pi^{\star}$ simultaneously. In particular, roughly $n=O\left(\frac{\mathcal{C}^{\pi^{\star}}\log(\lvert\Pi\rvert/\delta)}{\varepsilon^{2}}\right)$ examples are sufficient to learn a policy that is $\varepsilon$-suboptimal relative to $\pi^{\star}$. As a result, $\chi$PO is robust to overoptimization since the learned policy is as good as any $\pi^{\star}$ that is sufficiently covered by $\pi_{\mathsf{ref}}$ (in the sense that $\mathcal{C}^{\pi^{\star}}=O(1)$), which is effectively the best one can hope for in the purely offline setting. In contrast, naive offline alignment methods like DPO have sample complexity that scales with all-policy concentrability (roughly, $\max_{\pi}\mathcal{C}^{\pi}$), even when the comparator policy $\pi^{\star}$ is sufficiently covered (Zhu et al., 2023; Song et al., 2024).
[…]
Regarding the parameter $V_{\mathsf{max}}$, we observe that since the policy $\pi^{\star}_{{\scriptscriptstyle\beta}}$ satisfies $\left\lvert\beta\phi\left(\frac{\pi^{\star}_{{\scriptscriptstyle\beta}}(a\mid{}x)}{\pi_{\mathsf{ref}}(a\mid{}x)}\right)-\beta\phi\left(\frac{\pi^{\star}_{{\scriptscriptstyle\beta}}(b\mid{}x)}{\pi_{\mathsf{ref}}(b\mid{}x)}\right)\right\rvert\leq 2R_{\mathsf{max}}$, information-theoretically we can always achieve $V_{\mathsf{max}}=2R_{\mathsf{max}}$ by pre-filtering the policy class $\Pi$ to remove all policies for which this inequality does not hold. Since this may be non-trivial in practice, we incorporate clipping in Eq. 9, which, for precisely the above reason, we expect to improve performance empirically. In Section 4, we discuss the role of the $V_{\mathsf{max}}$ parameter and Assumption 3.2 in greater depth. See also the guarantees for the $\chi^{2}$-RLHF algorithm in Appendix B, which avoid dependence on this parameter.
@@
Consider the setting in Example 3.1, where the policy class $\Pi_{\mathcal{R},\beta}$ is the set of mixed $\chi^{2}$-regularized policies induced by a reward model class $\mathcal{R}$ with $r^{\star}\in\mathcal{R}$ and $\beta>0$. For any $\delta\in(0,1)$, there exists a choice [^5] It is unclear how to select $\beta^{\star}$ in a data-driven manner, as it depends on the (unknown) functionals $\pi\mapsto C^{\pi}$ and $\pi\mapsto J(\pi)$. for $\beta^{\star}>0$ such that with probability at least $1-\delta$, $\chi$PO (Algorithm 1), with class $\Pi_{\mathcal{R},\beta^{\star}}$, produces a policy $\widehat{\pi}$ such that for all policies $\pi^{\star}$ simultaneously, we have
@@
Specializing to the case of multi-armed bandits, we believe that the sample complexity bound in Eq. 13 is optimal in general (Rashidinejad et al., 2021). Note that while we consider finite classes $\Pi$ in Theorem 3.1 for simplicity, extension to infinite classes is trivial via standard uniform convergence arguments. We remark that the exponential dependence on $R_{\mathsf{max}}$ in Theorem 3.1 is an intrinsic feature of the Bradley-Terry model, and can be found in all prior work (Rosset et al., 2024; Xie et al., 2024). Finally, we remark that the weight on the KL term in the mixed $\chi^{2}$-regularized objective is not important for our statistical guarantees. For any $\gamma\in(0,1]$, we can replace the link function $\phi(\cdot)$ in $\chi$PO with $\phi_{\gamma}(z)=z+\gamma\log{}z$,
@@
which corresponds to the regularized objective $J_{\beta,\gamma}^{\chi_{\textsf{mix}}}(\pi)=\operatorname{\mathbb{E}}_{\pi}\left[r^{\star}(x,a)\right]-\beta\cdot D_{\chi^{2}}\left(\pi\;\|\;{}\pi_{\mathsf{ref}}\right)-\gamma\beta\cdot D_{\mathsf{KL}}\left(\pi\,\|\,\pi_{\mathsf{ref}}\right)$. This leads to identical guarantees for any $\gamma\in(0,1]$ (Theorem F.1 in Appendix F); essentially, we only require that $\gamma$ is positive to ensure the reparameterization in Eq. 10 is admissible.
@@ 4.2 The Bias-Overoptimization Tradeoff
Bias. Compared to DPO, $\chi$PO has a higher probability of taking both the optimal action $a_{\mathsf{good}}$ and the reference action $a_{0}$. As a result, it strikes a better bias-overoptimization tradeoff than DPO, and is competitive with respect to the comparator $a_{0}$ even when DPO fails to converge.
@@ 4.3 An Illustrative Example
Figure 3: The regret $J(a_{0})-J(\widehat{\pi})$ in the construction from Proposition A.1 for different values of $n$. We again condition on the “bad” event $\mathcal{E}$ where ${\widehat{r}}=r_{2}\neq r^{\star}$. For each $n$, the error from overoptimization dominates when $\beta\leq(2\log n)^{-1}$ (as discussed in Section 4.3), and the error from bias dominates when $\beta>(2\log n)^{-1}$. Taking the best choice of $\beta$ for each method, DPO converges at an exponentially slower rate than $\chi$PO.
@@ 4.4 Nontriviality and Role of $V_{\mathsf{max}}$ Parameter
Assumption 3.2 effectively implies that all policies $\pi\in\Pi$ satisfy $\big{\|}\frac{\pi}{\pi_{\mathsf{ref}}}\big{\|}_{\infty}\lesssim\frac{V_{\mathsf{max}}}{\beta}$; in other words, the policy class we use in $\chi$PO satisfies all-policy $L_{\infty}$-concentrability with $\max_{\pi\in\Pi}\mathcal{C}^{\pi}_{\infty}\lesssim\frac{V_{\mathsf{max}}}{\beta}$. At first glance, this might seem to trivialize the offline alignment problem, since it would suffice to prove a generalization guarantee based on all-policy concentrability, and then plug this bound in. We will show that this is not the case, and that this is actually an intrinsic feature of $\chi^{2}$-regularization.
@@
In this section, we sketch the proof of the main guarantee for $\chi$PO, Theorem 3.1, with the full proof deferred to Appendix F. A central object in the proof is the implicit reward model induced by the $\chi$PO policy $\widehat{\pi}$, which we define via
@@
As we will show, this reward model is a natural bridge between $\chi$PO and the corresponding mixed $\chi^{2}$-regularized RLHF objective in Section 3.1, and allows us to view $\chi$PO from a reward-based perspective. In particular, note that if we analogously define an induced reward model class $\mathcal{R}_{\Pi}\vcentcolon={}\{r(x,a)=\beta\phi{}\left(\frac{\pi(a\mid{}x)}{\pi_{\mathsf{ref}}(a\mid{}x)}\right):\pi\in\Pi\}$, then 3 of $\chi$PO can be viewed as performing maximum likelihood estimation over this class (in the sense of Eq. 3) under the Bradley-Terry model. Under Assumption 3.1, $\mathcal{R}_{\Pi}$ realizes the true reward function $r$ up to an action-independent shift. As a result, if we define $\Delta^{{r}}(x,a,b)\vcentcolon={}r(x,a)-r(x,b)$, then using a fairly standard generalization bound for maximum likelihood estimation (e.g., Wong and Shen (1995); Zhang (2006); de Geer (2000); see Lemma F.1), we can show that
@@
Lemma 5.1 (Informal version of Lemma F.3).
@@
Lemma 5.2 (Informal version of Lemma F.2).
@@
6 Experiments in Offline Language Model Alignment
We perform preliminary evaluations of $\chi$PO for offline language model alignment on the TL;DR dataset (Stiennon et al., 2020), using DPO as our comparison baseline. The reference policy $\pi_{\mathsf{ref}}$ is the Pythia-1b model (Biderman et al., 2023) pre-trained on SFT data (cleanrl/EleutherAI_pythia-1b-deduped__sft__tldr from Huang et al. (2022)), and performance is measured via winrate against a baseline, as judged by GPT-4o. All parameters that are not algorithm-specific, such as the learning rate, are shared by both $\chi$PO and DPO in order to ensure a fair comparison (see Appendix C for details).
Table 1: Winrate on TL;DR Summarization for models learned by $\chi$PO and DPO, for several choices of the number of training epochs and the regularization parameter $\beta$. Standard error over 3 seeds is also reported.
$\beta$ Epochs $\chi$PO winrate (%) DPO winrate (%)
0.05 1 $\mathbf{56.5\pm 1.3}$ $55.8\pm 2.1$
2 $\mathbf{56.1\pm 0.6}$ $50.3\pm 0.8$
4 $\mathbf{48.0\pm 1.6}$ $38.0\pm 0.7$
0.005 1 $\mathbf{50.6\pm 1.6}$ $14.7\pm 3.9$
2 $\mathbf{52.8\pm 2.3}$ $3.4\pm 1.5$
4 $\mathbf{51.6\pm 0.8}$ $0.5\pm 0.2$
In Table 1 we display the winrates of $\chi$PO and DPO over several choices of training epochs, as well as regularization parameter $\beta$. The winrate corresponds to the final checkpoint learned by each algorithm for each set of hyperparameters. We consider $\beta=0.05$ and 1 epoch of training to be a standard setup for DPO (Gao et al., 2024; Guo et al., 2024; Rafailov et al., 2024a), and, as we are particularly concerned with regimes where overoptimization is of concern, we additionally analyze performance when epochs are increased, and/or $\beta$ is decreased (corresponding to less regularization).
Figure 4: (Left) TL;DR Summarization winrate recorded longitudinally over 2 epochs of training every 250 steps. Shaded area displays ±$\pm 1$ standard error over 3 seeds. At 1 epoch $\chi$PO already obtains better performance, and continues to improve over the course of training, while DPO degrades over time. (Right) KL divergence $D_{\mathsf{KL}}\left(\widehat{\pi}\,\|\,\pi_{\mathsf{ref}}\right)$ averaged over 2 of the seeds. For the same $\beta$, $\chi$PO constrains the learned policy to be significantly closer to $\pi_{\mathsf{ref}}$, thereby striking a better bias-variance tradeoff.
Over all choices of $\beta$ and epochs, $\chi$PO achieves a higher average winrate than DPO. While the difference is not significant for $\beta=0.05$ and 1 epoch, the performance gap grows significantly as the number of epochs increases, demonstrating the robustness of $\chi$PO to overoptimization. Further, while DPO degrades completely for $\beta=0.005$, $\chi$PO is robust over two orders of magnitude of $\beta$, reinforcing trends seen earlier in Fig. 3 and the more favorable bias-overoptimization tradeoff from our theoretical analysis.
In addition, $\chi$PO exhibits better performance and robustness longitudinally throughout training, as shown in Fig. 4. While DPO peaks early with high variance around 0.5 epochs and degrades thereafter, $\chi$PO continues to improve smoothly then plateaus over the last epoch. Further, for the same regularization parameter $\beta$, the $\chi$PO policy has significantly lower KL-divergence relative to $\pi_{\mathsf{ref}}$, demonstrating that the $\chi^{2}$-regularization is both a stronger regularizer and one that effectively mitigates overoptimization.
7 $\boldsymbol{\chi}$PO for General Preference Models